OpenStudy (anonymous):

integral of dx/x^p from x=o to x=1

OpenStudy (anonymous):

\[ \int_0^1 \dfrac {dx}{x^p} = \int_0^1 x^-p \]Do you know how to do the indefinite integral. This is just reverse differentiation.

OpenStudy (anonymous):

you can re-write it as (1/x^P)dx, and further re-write it as x^-p dx. you should be able to do it from there

OpenStudy (anonymous):

Did you get your answer? The two above give correct answer. I'll just rewrite it :) \[\int_0^1\frac {dx}{x^p}=\int_0^1x^{-p}dx\]

OpenStudy (anonymous):

Oops! Thanks for the correcting my LaTeX. That's what I meant!